1+2+3=154 2+3+4=276 3+4+5=398 4+5+6=???????????????????????????? ONLY FOR GENIOUS.................................................. 99 % FAIL

1+2+3=154
2+3+4=276
3+4+5=398
4+5+6=????????????????????????????
ONLY FOR GENIOUS..................................................
99 %  FAIL

Grade:11

2 Answers

Tanuja
6 years ago
According to the given series the answer is 41110 I guess is it correct 4 5+6 6+4 on simplifying it is 41110
venkata ramana
6 years ago
hi naveen
1+2+3=154
in 154 first digit comes from given first letter 1
            second digit comes from given sum of second and third letters 5
            third digit comes from given sum of first and third letters 4
2+3+4=276
it is also same process
2,sum of 3and4,sum of 2and 4
 
SAME PROCESS 4+5+6=41110